Five years ago, you purchased 220.80 shares of a mutual fund. Since then, you have reinvested your fund dividends and acquired an additional 35.10 shares. The fund currently has an NAV of $37.60. The fund charges a contingent deferred sales charge of 5 percent for the first 2 years after which time the charge declines by 1 percent a year. How much money will you receive if you redeem all of your shares today?

Group of answer choices

$9,700.79

$9,510.18

$9,429.40

$9,360.32

$9,633.75

2.

Five months ago, you purchased 200 shares of a mutual fund at an offering price of $54 a share. The fund imposes a front-end load of 4.5 percent and has total annual expenses of 1.08 percent. The NAV of the fund today is $52.40. There were no fund distributions during these five months. What is your holding period return on this investment?

Group of answer choices

1.89 percent

-1.92 percent

-2.96 percent

2.26 percent

1.44 percent
Answer #1

1)

Total number of shares = 220.80+35.10 = 255.90

NAV = $37.60

charges for first two years = 5%

3rd year = 4%(because it declines one percent per year from year 2 on wards)

4th year = 3%

5th year = 2%

total money to be received = 255.90 x 37.60 x (1 - 0.02) = 9429.40

2)

Holding period return = (present NAV - purchase price) / purchase price

= (52.40 - 54) / 54

= -2.96%